OpenStudy (anonymous):

ln e = ln sqrt 7/x-ln e find the exact solution.

OpenStudy (anonymous):

lne=1 1=ln(sqrt(7)/x)-1

OpenStudy (anonymous):

so how do i figure out what x equals?

OpenStudy (anonymous):

\[\ln e =\ln \frac{ \sqrt{7} }{ x }-\ln e\] this is the equation that I have to find the exact solution for. I'm sorry I didn't know I could make the equation look like this.

OpenStudy (agent0smith):

Okay, then you can use some log laws to simplify it: log(a/b) = log a - log b \[1 =\ln \frac{ \sqrt{7} }{ x }-1\] \[1 =\ln \sqrt{7} - \ln{ x }-1\] \[\ln x = \ln \sqrt7 - 2\] Then take both sides to the power of e.